OpenStudy (pinkgirl56):

lim (x)/(x+1) x -> 1- Whenever i do this I keep getting 1/2 but the choices are : a. -inf b. 0 c. -1 d. inf

OpenStudy (irishboy123):

yep, it's continuous through that limit: \(\dfrac{1}{1 + 1}\) so check for typo's ....dunno

OpenStudy (pinkgirl56):

@IrishBoy123

OpenStudy (pinkgirl56):

the answer was infinity

OpenStudy (irishboy123):

cos you typed out the wrong question?? \(\large \lim\limits_{x \to -1}\) might be what you meant \(\large \lim\limits_{x \to 1^{-}}\) is what you typed